The daily revenue of Modern Style Hair Salon is approximately normally distributed with a mean of R7200 and a standard deviation of R1500. 75% of days, the salon makes more than? (Rands only no cents)


1
Expert's answer
2021-11-15T19:08:15-0500
"P(X>x)=1-P(Z\leq\dfrac{x-7200}{1500})=0.75"

"\dfrac{x-7200}{1500}=-0.674490"

"x=7200-0.674490(1500)"

"x=R6188"